what is the number of sides of a regular polygon whose interior angles measure 135 degrees?
Answer by Alan3354(69443)

what is the number of sides of a regular polygon whose interior angles measure 135 degrees?
--------
Ext angles = 180 - 135 = 45
n = 360/45 = 8 sides